OpenStudy (anonymous):

Whats the derivative of e^(x^1/2)

OpenStudy (anonymous):

using chain rule you take the derivative of the outer function, keep the inner function and the multiply it by the derivative of the inner function so derivative of outer = e^(x^(1/2)) derivative of innner = 1/2x^(-1/2) so your answer is those 2 multiplied together

OpenStudy (anonymous):

\[\begin{array}{rcl} y = e^{\sqrt{x}} \\ \ln(y)=\sqrt{x} \end{array} \]

OpenStudy (anonymous):

\[ (\ln(y))' = \frac{y'}{y} \]

OpenStudy (anonymous):

\[ (\sqrt{x})'=(x^{1/2})'=\frac{1}{2}x^{-1/2} = \frac{1}{2\sqrt{x}} \]

OpenStudy (anonymous):

thank u so much!

OpenStudy (anonymous):

np glad to help